OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____. a. 3x = 2 b. 3x = 5 c. 5x = 5

OpenStudy (dmndlife24):

2(2x + y = 3) 4x + 2y = 6 x - 2y = -1 y values cancel out... 5x = 5 divide both sides by 5... x = 1

OpenStudy (anonymous):

thank you!